OpenStudy (anonymous):

How do you simplify this:(k^7)^3

OpenStudy (anonymous):

power to a power I'm assuming?

OpenStudy (turingtest):

\[\huge(k^7)^3\]like so?

OpenStudy (anonymous):

if they are both powers, you just multiply them together.

OpenStudy (anonymous):

Yeah thats how i tried to put, sorry my computer typed it funny

OpenStudy (anonymous):

Yeah you just multiply them together

OpenStudy (turingtest):

the formula @Allie13 is referring to is\[\huge(x^a)^b=x^{ab}\]as she said, you multiply the exponents in cases like this

OpenStudy (anonymous):

(k^7)^3= k^(3*7) = k^(21)

OpenStudy (anonymous):

^all right :D

OpenStudy (turingtest):

can everyone please stop typing the final answer? I think Allie and I have provided enough info to solve it.

OpenStudy (anonymous):

Dont worry 'bout it i get it now, thanks :)

OpenStudy (turingtest):

welcome!